OpenStudy (campayne):

Use elimination to solve the system of equations. Enter your answer as an ordered triple; example: (2, -3, 4) {4x + y + 3z = 0 2x - 2y - z = 10 3x - 2y + 2z = 11

OpenStudy (jango_in_dtown):

Gaussian elimination?

OpenStudy (campayne):

ive not in the slightest idea

OpenStudy (callielovesyhuu):

We have the following system: 4x+y+3z=0 2x-2y-z=10 3x-2y+2z=11 so, we have the following matrix. 4 1 3 0 2 -2 -1 10 3 -2 2 11 We start: 4 1 3 0 0 -5 -5 20 2R₂-R₁ 0 -11 -1 44 4R₃-3R₁ 4 1 3 0 0 -5 -5 20 0 0 50 0 5R₃-11R₂ Then: 50Z=0 ⇒ Z=0/50=0 -5y-5Z=20 -5y-5(0)=20 -5y-0=20 -5y=20 y=-20/5 y=-4 4x+y+3z=0 4x-4+3(0)=0 4x-4=0 4x=4 x=4/4 x=1 Answer: (1,-4,0)